LMNP is a rectangle. LN= 9x-14 and MP=7x+4.What is the length of the diagonals?
Answer by nyc_function(2741)

9x-14 = 7x + 4
9x-7x = 4 + 14
2x = 18
x = 9
9(9) -4 = 81 minus 4 = 77
7(9)+4 = 63 + 4 = 67
Then (77)^2 + (67)^2 = (hypotenuse)^2
Can you take it from here?